# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ 5.3 Implications for measuring probabilistic extraction on books
Three important observations follow from these results (and related experiments in Appendix E.4); we highlight them here and defer additional discussion to Appendix E. First, the negative controls provide strong evidence that our sliding-window probabilistic extraction procedure is valid. Second, they indicate that our choice of $\tau_{\text{min}}$ for counting extraction success is indeed very conservative. Given the very low probabilities we observe for generating non-training data (Figure 7, Appendix E.4.2), we could reasonably set $\tau_{\text{min}}$ below $0.1\%$. As such, our results almost certainly under-count memorization, and should be read as a lower bound. Nevertheless, we prefer to err on the side of caution. To lower $\tau_{\text{min}}$ and remain confident about validity, we would need to run a much larger set of negative controls, which exceeds the compute budget for this work. Part of the reason is that, as others have noted [28], LLMs are not actually like monkeys randomly outputting tokens. They have learned to generate structured, grammatical sentences, so the distribution of plausible outputs is far smaller than random chance would suggest (Appendix E.5). Exploring what this implies for how high $p_{\bm{z}}$ must be to count as evidence of memorization is an important question for future work. For now, we prefer to under-count memorization rather than risk over-counting it.
@@ 6.1 Experimental setup
We use the same setup as in Section 4.1: the sliding-window procedure with top-$k$ decoding (T=1T{=}1, k=40k{=}40) for $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). We run this procedure on $17$ models, which we organize into three groups: (a) models trained on Books3 (Llama [127, 128, 54] and Pythia [9], blue); (b) models that were very likely trained on (at least parts of) books also contained in Books3, based on measurements of $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$ for unique content from those books (DeepSeek v1, Qwen 2.5, and Gemma 2, red); and (c) a model not trained on copyrighted books (Phi 4, orange). Of the ${\sim}200{,}000$ books in Books3, we include $50$ in our analysis. These are the same books for which we analyze extraction coverage in Figure 5. They include works associated with plaintiffs in the amended class action complaint of Kadrey et al. v. Meta, Inc. [70, pp. 4-5], [^12] That suit was recently decided in favor of the defendants (Meta) [98]. as well as generally popular books [e.g., 106, 126, 13, 61]), and some academic texts [e.g., 139, 7]). We suspected that some popular books would exhibit higher degrees of memorization (e.g., due to duplicated text from other sources). This is why we did not select books at random; we deliberately chose titles intended to capture variation across Books3. Further details on these decisions are in Appendix D.1. We ran thousands of experiments on these $50$ books and $17$ models (Appendix D.3). (In total, these experiments cost $6$ GPU months.) Nevertheless, our results should not be read as a complete account of memorization across the entire Books3 dataset.
@@ 6.2 Key patterns in our results
Comparing a single book across models.  Each book in Figure 8 illustrates a different high-level pattern of memorization across models. There are books like M. Butterfly [65], for which our sliding-window procedure reveals almost no memorization: aside from the copyright notice, we extract little—if any—text from these books across models of different sizes trained by different organizations. Most books we test with $50$-token prefixes resemble this pattern. However, for a given book, we are often able to extract some substantive sequences from Llama 3 and 3.1 70B.
@@
Books like Beloved [89] represent a more intermediate case: for these books, memorization is generally low for most models with respect to $50$-token prefixes (Appendix E.3), but with important exceptions. Llama 3 and 3.1 70B memorize a considerable number of fragments in the book. Other Llama models memorize significantly less, and category (b) models (DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B) memorize even less. For Beloved, the $7$ largest models we test—category (b) and all large Llama models—memorize the same passages at ${\sim}170$[90, 92] and ${\sim}480$K characters [91]. Both are famous quotes from the book, which we confirmed are highly duplicated on the Internet. As with Phi 4 (Section 5.2), models that otherwise exhibit low memorization (except for such passages) may not have been trained on the whole book; rather, they may have memorized these passages due to their inclusion and duplication in other sources (Appendix E.6).
@@
Last, select books exhibit enormous degrees of memorization for the $7$ largest models, with lesser amounts of memorization in smaller models like Llama 1 13B. These are widely popular books like Harry Potter and the Sorcerer’s Stone [106] (Figure 8) and 1984 [100] (Figure 3). While memorization of such books tends to be significant for larger models, it is especially high for Llama 3 and 3.1 70B, which appear to memorize several books in their entirety [e.g., 107, 46, 68, 126].
@@
Comparing different books within a model.  Small models generally memorize less than larger ones, but there is variation across books. For example, Llama 1 13B memorizes more of Harry Potter and the Sorcerer’s Stone than M. Butterfly and Beloved; Llama 3.1 8B memorizes portions of We Were Eight Years in Power [23] (Figure 6), but less from other books (Appendix E.3). Variation is more pronounced in larger models. As noted above, for all models, most results resemble those shown in Figure 8 for M. Butterfly. This is true for most of the plaintiff-authored books in Kadrey et al. v. Meta, Inc. [70]—with exceptions (e.g., Llama 3.1 70B on Coates [23]; see Figure 6). Extraction coverage distributions reinforce this conclusion: they skew toward low coverage (Section 4.2, Appendix D.4). However, there are clearly several exceptions. Llama 3 and Llama 3.1 70B memorize significantly more from the $50$ books than any other model we tested. Many of these books yield results similar to those for Beloved [e.g., 5, 11, 13, 57]. In extreme cases, both models appear to memorize entire books, such as Harry Potter and the Sorcerer’s Stone. We address this further in Section 7. Llama 1 65B, Llama 2 70B, and category (b) models also exhibit high memorization for some, but not all, of these books.
@@
Altogether, these patterns underscore the importance of training choices in shaping memorization. Models do not memorize all of their training data [74, 16, 84, 26]. The mere existence of particular sequences within a training dataset does not imply a particular conclusion about memorization [74, 75, 76, 77]. For example, similarly sized models trained by different organizations display different degrees of memorization of the same book—even when the evidence strongly suggests the entire book (e.g., Harry Potter and the Sorcerer’s Stone) was included in their respective training datasets (Appendix E.6). This is also evident with respect to just Llama models, which were all trained by the same organization (Meta) on Books3. All Llama models (by size) memorize more than the other models we evaluate. But notably, within size classes, newer versions typically memorize more from specific books than earlier ones: Llama 3 models appear to have memorized more from books than Llama 2, which in turn memorize more than Llama 1.
@@ 7.1 A simple reconstruction procedure
With appropriate choices of $s$, $n$, $m$, and configuration of beam search, [^14] Beam search is a deterministic decoding method that explores multiple candidate sequences in parallel to approximate the highest-probability continuation under the model. At each generation step, it keeps the top-$k$ highest-probability sequences so far. At the final step, it outputs the sequence with the highest probability. we attempt to generate the entire book with high fidelity. On our hardware, after testing just a handful of configurations for beam search and a few different short seed prompts, we were able to deterministically generate a near-exact copy of the entire 304-page book. We show results in this section using a seed prompt of just "Mr. and Mrs. D" (s=7s{=}7 tokens, drawn from the first line of the first chapter) and a beam width of k=8k{=}8; we use a sliding context length of n=3000n{=}3000 tokens and generate m=50m{=}50 tokens at each step. In practice, the only minor complication was that Llama 3.1 70B tended to predict end-of-sequence (EOS) tokens at the ends of chapters. To address this, we removed the EOS token from the generated $m$ tokens before appending them to the context, and replaced it with the tokens for "CHAPTER {c+1}" (with {c+1} spelled out in words, e.g., when c=2c{=}2, we insert "CHAPTER THREE"). With this simple adjustment, the model successfully continued to generate the book.
@@
Figure 9: Samples of the diff between the ground-truth text of Harry Potter and the Sorcerer’s Stone [106] from Books3 and the entire book text we generated using Llama 3.1 70B, starting with a short seed prompt from the first chapter. See Appendix F.1.
@@ 8.1 LLMs could be derivative works of the books they memorize
In the case of some models, there is quite a lot of memorization of some books, though most books are not memorized at all—with respect to our specific extraction methodology—either in whole or in part. Such evidence of memorization in the model matters for the dispute over whether models themselves are derivative works (Section 3), an argument that courts have thus far not been receptive to [99, 98].
@@
A work is not a derivative work unless it is “substantially similar” in significant part to the original work [e.g., 82]. While prior debates over whether models can qualify as derivative works have been largely theoretical [28, 76], the results of our probabilistic extraction procedure offer empirical evidence that is directly relevant to this issue. They indicate that models are not, as plaintiffs sometimes contend, mere copies of all of the works on which they were trained. However, they also suggest that at least some LLMs (e.g., Llama 3.1 70B) may be derivative works of at least some books (e.g., Harry Potter and the Sorcerer’s Stone)—because those LLMs have memorized a significant amount of protectable expression from those books. The law does not require that the entire work be included in the derivative; it is enough that the derivative incorporates a substantial amount of protectable expression. We find that this turns out to be true for some (but not all) books for some (but not all) models.
@@
Fair use.  The inquiry does not end here; the model, like the training dataset, may be protected under copyright’s fair use doctrine [79]. But the fair use analysis of the model itself may be different than the fair use analysis of the training dataset [76]. Copyright law treats intermediate copies, such as those used internally within a system, differently from external-facing copies. A sequence of copyrighted data used for training is an intermediate copy used internally by AI companies for model training, rather than sold as a product. Unless that sequence is extracted, it is only used internally in the course of the model producing outputs. Beyond extraction, individual sequences may to lesser or greater extents influence model outputs, but many of these outputs will not be substantially similar to any training data [62, 76].
@@
A model, too, is an intermediate work in many cases. If the model memorizes content but only generates noninfringing outputs, that is a strong argument that how the model makes use of that content is transformative and ought to be permissible fair use. But for companies that sell or release their models to others under open source licenses, the model itself is the product, and sometimes one that is being sold for commercial gain. That does not defeat the fair use argument, but it makes the analogy to training data less precise. In these instances, it may be harder to rely on the cases that justify training on copyrighted data as fair use [e.g., 6, 115, 120].
@@
Potential consequences of a model being a copy of works it has memorized.  A finding that the model itself is a copy of some copyrighted works, and that a distribution of the model was thus a distribution of a copy of those works, could be dramatic for the AI industry. Llama 3.1 70B was downloaded 105,029105{,}029 times from HuggingFace in May 2025 (Appendix G)—far less from the height of its release in 2024. If we say conservatively that the model has been downloaded 1 million times since its release, [^16] Meta announced in March 2025 that total downloads for all of its Llama models surpassed $1$ billion [86]. In 2024, the Times reported that Llama 2 models had been downloaded $180$ million times [66]. then those 1 million downloads could be seen as 1 million potentially infringing distributions of reproductions of Harry Potter and the Sorcerer’s Stone—as well as any other books in copyright for which a more than de minimis amount of copyrighted expression has been memorized.
@@
In general, damages for copyright infringement can run very high: statutory damages can be as high as \$150,000 per infringed work, and actual damages can be much higher. If the model itself is a copy and is not protected by fair use, courts might also order that it not be distributed or even that copies of the model itself be destroyed [31]. We have only studied a relatively small number of books ($50$) and models ($17$), so cannot draw strong conclusions about the extent of memorization of books in general. In principle, if the overall quantity of memorized books in open-weight models is particularly high, it is not an exaggeration to say that the risk of damages across all of those books and of orders of destruction are an existential threat to the companies training these models.
@@
Policy considerations.  The effect of these copyright rules may be to treat openly distributed models less favorably than models gated behind closed system, because the distributed models contain a latent copy or derivative of some copyrighted works. However, there is no good policy reason to treat open-weight models more harshly in copyright law than proprietary ones. Policymakers may want to act to protect models from being declared unlawful copies, in order to prevent crushing financial liability on developers for providing a model that generates transformative, noninfringing outputs. However, a blanket protection for all open-weight models may also be unwarranted. The extent to which models memorize their training data is significantly influenced by choices during training, such as de-duplicating the training dataset [28, 18, 76, 74, 60] Policymakers may be less inclined to protect models if the developers made training choices known to substantially increase the risk of memorization—for example, deliberately duplicating particular books.

### v4

@@ 5.3 Implications for measuring probabilistic extraction on books
Three important observations follow from these results (and related experiments in Appendix E.4); we highlight them here and defer additional discussion to Appendix E. First, the negative controls provide strong evidence that our sliding-window probabilistic extraction procedure is valid. Second, they indicate that our choice of $\tau_{\text{min}}$ for counting extraction success is indeed very conservative. Given the very low probabilities we observe for generating non-training data (Figure 7, Appendix E.4.2), we could reasonably set $\tau_{\text{min}}$ below $0.1\%$. As such, our results almost certainly under-count memorization, and should be read as a lower bound. Nevertheless, we prefer to err on the side of caution. To lower $\tau_{\text{min}}$ and remain confident about validity, we would need to run a much larger set of negative controls, which exceeds the compute budget for this work. Part of the reason is that, as others have noted [cooper2024files], LLMs are not actually like monkeys randomly outputting tokens. They have learned to generate structured, grammatical sentences, so the distribution of plausible outputs is far smaller than random chance would suggest (Appendix E.5). Exploring what this implies for how high $p_{\bm{z}}$ must be to count as evidence of memorization is an important question for future work. For now, we prefer to under-count memorization rather than risk over-counting it.
@@ 6.1 Experimental setup
We use the same setup as in Section 4.1: the sliding-window procedure with top-$k$ decoding (T=1T{=}1, k=40k{=}40) for $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). We run this procedure on $17$ models, which we organize into three groups: (a) models trained on Books3 (Llama [touvron2023llamaopenefficientfoundation, llama2, llama3] and Pythia [biderman2023pythia], blue); (b) models that were very likely trained on (at least parts of) books also contained in Books3, based on measurements of $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$ for unique content from those books (DeepSeek v1, Qwen 2.5, and Gemma 2, red); and (c) a model not trained on copyrighted books (Phi 4, orange). Of the ${\sim}200{,}000$ books in Books3, we include $50$ in our analysis. These are the same books for which we analyze extraction coverage in Figure 5. They include works associated with plaintiffs in the amended class action complaint of Kadrey et al. v. Meta, Inc. [kadreyamendedconsolidated, pp. 4-5], [^12] That suit was recently decided in favor of the defendants (Meta) [kadreyjudgment]. as well as generally popular books [e.g., hp1, hobbit, sisyphus, catch22]), and some academic texts [e.g., zittrain, dante]). We suspected that some popular books would exhibit higher degrees of memorization (e.g., due to duplicated text from other sources). This is why we did not select books at random; we deliberately chose titles intended to capture variation across Books3. Further details on these decisions are in Appendix D.1. We ran thousands of experiments on these $50$ books and $17$ models (Appendix D.3). (In total, these experiments cost $6$ GPU months.) Nevertheless, our results should not be read as a complete account of memorization across the entire Books3 dataset.
@@ 6.2 Key patterns in our results
Comparing a single book across models.  Each book in Figure 8 illustrates a different high-level pattern of memorization across models. There are books like M. Butterfly [mbutterfly], for which our sliding-window procedure reveals almost no memorization: aside from the copyright notice, we extract little—if any—text from these books across models of different sizes trained by different organizations. Most books we test with $50$-token prefixes resemble this pattern. However, for a given book, we are often able to extract some substantive sequences from Llama 3 and 3.1 70B.
@@
Books like Beloved [beloved] represent a more intermediate case: for these books, memorization is generally low for most models with respect to $50$-token prefixes (Appendix E.3), but with important exceptions. Llama 3 and 3.1 70B memorize a considerable number of fragments in the book. Other Llama models memorize significantly less, and category (b) models (DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B) memorize even less. For Beloved, the $7$ largest models we test—category (b) and all large Llama models—memorize the same passages at ${\sim}170$K [belovedquote1, belovedquote3] and ${\sim}480$K characters [belovedquote2]. Both are famous quotes from the book, which we confirmed are highly duplicated on the Internet. As with Phi 4 (Section 5.2), models that otherwise exhibit low memorization (except for such passages) may not have been trained on the whole book; rather, they may have memorized these passages due to their inclusion and duplication in other sources (Appendix E.6).
@@
Last, select books exhibit enormous degrees of memorization for the $7$ largest models, with lesser amounts of memorization in smaller models like Llama 1 13B. These are widely popular books like Harry Potter and the Sorcerer’s Stone [hp1] (Figure 8) and 1984 [1984] (Figure 3). While memorization of such books tends to be significant for larger models, it is especially high for Llama 3 and 3.1 70B, which appear to memorize several books in their entirety [e.g., hp4, gatsby, ulysses, hobbit].
@@
Comparing different books within a model.  Small models generally memorize less than larger ones, but there is variation across books. For example, Llama 1 13B memorizes more of Harry Potter and the Sorcerer’s Stone than M. Butterfly and Beloved; Llama 3.1 8B memorizes portions of We Were Eight Years in Power [wewere] (Figure 6), but less from other books (Appendix E.3). Variation is more pronounced in larger models. As noted above, for all models, most results resemble those shown in Figure 8 for M. Butterfly. This is true for most of the plaintiff-authored books in Kadrey et al. v. Meta, Inc. [kadreyamendedconsolidated]—with exceptions (e.g., Llama 3.1 70B on wewere; see Figure 6). Extraction coverage distributions reinforce this conclusion: they skew toward low coverage (Section 4.2, Appendix D.4). However, there are clearly several exceptions. Llama 3 and Llama 3.1 70B memorize significantly more from the $50$ books than any other model we tested. Many of these books yield results similar to those for Beloved [e.g., handmaid, davinci, sisyphus, curious]. In extreme cases, both models appear to memorize entire books, such as Harry Potter and the Sorcerer’s Stone. We address this further in Section 7. Llama 1 65B, Llama 2 70B, and category (b) models also exhibit high memorization for some, but not all, of these books.
@@
Altogether, these patterns underscore the importance of training choices in shaping memorization. Models do not memorize all of their training data [lee2022dedup, carlini2021extracting, mahdavi2024memorizationcapacitymultiheadattention, collins2017capacitytrainabilityrecurrentneural]. The mere existence of particular sequences within a training dataset does not imply a particular conclusion about memorization [lee2022dedup, lee2023explainers, lee2023talkin, lee2024talkinshort]. For example, similarly sized models trained by different organizations display different degrees of memorization of the same book—even when the evidence strongly suggests the entire book (e.g., Harry Potter and the Sorcerer’s Stone) was included in their respective training datasets (Appendix E.6). This is also evident with respect to just Llama models, which were all trained by the same organization (Meta) on Books3. All Llama models (by size) memorize more than the other models we evaluate. But notably, within size classes, newer versions typically memorize more from specific books than earlier ones: Llama 3 models appear to have memorized more from books than Llama 2, which in turn memorize more than Llama 1.
@@ 7.1 A simple reconstruction procedure
With appropriate choices of $s$, $n$, $m$, and configuration of beam search, [^14] Beam search is a deterministic decoding method that explores multiple candidate sequences in parallel to approximate the highest-probability continuation under the model. At each generation step, it keeps the top-$k$ highest-probability sequences so far. At the final step, it outputs the sequence with the highest probability. we attempt to generate the entire book with high fidelity. On our hardware, after testing just a handful of configurations for beam search and a few different short seed prompts, we were able to deterministically generate a near-exact copy of the entire 304-page book. We show results in this section using a seed prompt of just "Mr. and Mrs. D" (s=6s{=}6 tokens, drawn from the first line of the first chapter, prepended with the beginning-of-sequence token) and a beam width of k=8k{=}8; we use a sliding context length of n=3000n{=}3000 tokens and generate m=50m{=}50 tokens at each step. In practice, the only minor complication was that Llama 3.1 70B tended to predict end-of-sequence (EOS) tokens at the ends of chapters. To address this, we removed the EOS token from the generated $m$ tokens before appending them to the context, and replaced it with the tokens for "CHAPTER {c+1}" (with {c+1} spelled out in words, e.g., when c=2c{=}2, we insert "CHAPTER THREE"). With this simple adjustment, the model successfully continued to generate the book.
@@
Figure 9: Samples of the diff between the ground-truth text of Harry Potter and the Sorcerer’s Stone [hp1] from Books3 and the entire book text we generated using Llama 3.1 70B, starting with a short seed prompt from the first chapter. See Appendix F.1.
@@ 8.1 LLMs could be derivative works of the books they memorize
In the case of some models, there is quite a lot of memorization of some books, though most books are not memorized at all—with respect to our specific extraction methodology—either in whole or in part. Such evidence of memorization in the model matters for the dispute over whether models themselves are derivative works (Section 3), an argument that courts have thus far not been receptive to [bartzjudgment, kadreyjudgment].
@@
A work is not a derivative work unless it is “substantially similar” in significant part to the original work [e.g., litchfield]. While prior debates over whether models can qualify as derivative works have been largely theoretical [cooper2024files, lee2023talkin], the results of our probabilistic extraction procedure offer empirical evidence that is directly relevant to this issue. They indicate that models are not, as plaintiffs sometimes contend, mere copies of all of the works on which they were trained. However, they also suggest that at least some LLMs (e.g., Llama 3.1 70B) may be derivative works of at least some books (e.g., Harry Potter and the Sorcerer’s Stone)—because those LLMs have memorized a significant amount of protectable expression from those books. The law does not require that the entire work be included in the derivative; it is enough that the derivative incorporates a substantial amount of protectable expression. We find that this turns out to be true for some (but not all) books for some (but not all) models.
@@
Fair use.  The inquiry does not end here; the model, like the training dataset, may be protected under copyright’s fair use doctrine [lemley2021fairlearning]. But the fair use analysis of the model itself may be different than the fair use analysis of the training dataset [lee2023talkin]. Copyright law treats intermediate copies, such as those used internally within a system, differently from external-facing copies. A sequence of copyrighted data used for training is an intermediate copy used internally by AI companies for model training, rather than sold as a product. Unless that sequence is extracted, it is only used internally in the course of the model producing outputs. Beyond extraction, individual sequences may to lesser or greater extents influence model outputs, but many of these outputs will not be substantially similar to any training data [henderson2023fair, lee2023talkin].
@@
A model, too, is an intermediate work in many cases. If the model memorizes content but only generates noninfringing outputs, that is a strong argument that how the model makes use of that content is transformative and ought to be permissible fair use. But for companies that sell or release their models to others under open source licenses, the model itself is the product, and sometimes one that is being sold for commercial gain. That does not defeat the fair use argument, but it makes the analogy to training data less precise. In these instances, it may be harder to rely on the cases that justify training on copyrighted data as fair use [e.g., authorsguild, sega, sonyconnectix].
@@
Potential consequences of a model being a copy of works it has memorized.  A finding that the model itself is a copy of some copyrighted works, and that a distribution of the model was thus a distribution of a copy of those works, could be dramatic for the AI industry. Llama 3.1 70B was downloaded 105,029105{,}029 times from HuggingFace in May 2025 (Appendix G)—far less from the height of its release in 2024. If we say conservatively that the model has been downloaded 1 million times since its release, [^16] Meta announced in March 2025 that total downloads for all of its Llama models surpassed $1$ billion [metadownloads]. In 2024, the Times reported that Llama 2 models had been downloaded $180$ million times [nytllama2]. then those 1 million downloads could be seen as 1 million potentially infringing distributions of reproductions of Harry Potter and the Sorcerer’s Stone—as well as any other books in copyright for which a more than de minimis amount of copyrighted expression has been memorized.
@@
In general, damages for copyright infringement can run very high: statutory damages can be as high as \$150,000 per infringed work, and actual damages can be much higher. If the model itself is a copy and is not protected by fair use, courts might also order that it not be distributed or even that copies of the model itself be destroyed [17usc503]. We have only studied a relatively small number of books ($50$) and models ($17$), so cannot draw strong conclusions about the extent of memorization of books in general. In principle, if the overall quantity of memorized books in open-weight models is particularly high, it is not an exaggeration to say that the risk of damages across all of those books and of orders of destruction are an existential threat to the companies training these models.
@@
Policy considerations.  The effect of these copyright rules may be to treat openly distributed models less favorably than models gated behind closed system, because the distributed models contain a latent copy or derivative of some copyrighted works. However, there is no good policy reason to treat open-weight models more harshly in copyright law than proprietary ones. Policymakers may want to act to protect models from being declared unlawful copies, in order to prevent crushing financial liability on developers for providing a model that generates transformative, noninfringing outputs. However, a blanket protection for all open-weight models may also be unwarranted. The extent to which models memorize their training data is significantly influenced by choices during training, such as de-duplicating the training dataset [cooper2024files, carlini2023quantifying, lee2023talkin, lee2022dedup, hayes2025measuringmemorizationlanguagemodels] Policymakers may be less inclined to protect models if the developers made training choices known to substantially increase the risk of memorization—for example, deliberately duplicating particular books.